# On the maximal correlation of some stochastic processes

### v2

@@ 5.6 Examples
Firstly, we consider an example studied in [BS25]. They have calculated the maximal correlation for the bivariate Marshall Olkin exponential distribution. We state their result in the following theorem:
@@
Let $W_{1},W_{2}$ and $W_{3}$ be independent exponential random variables. For $i=1,2,3$, the parameter of $W_{i}$ is $\lambda_{i}>0$. Let $V_{1}=\min(W_{1},W_{3})$ and $V_{2}=\min(W_{2},W_{3})$. Then, we have that
@@
by using Corollary 1.8 and Lemma 2.3.
We take i.i.d. random variables $X_{1},X_{2},\ldots,X_{n}$. Each $X_{i}$ is an exponential random variable with parameter $\lambda$. Let $(S,T)$ be the pair of random subsets of $[n]$ defined in Corollary 1.8. Let $W_{1}=\min_{i\in S\setminus T}X_{i}$, $W_{2}=\min_{i\in T\setminus S}X_{i}$ and $W_{3}=\min_{i\in S\cap T}X_{i}$. Since $(S,T)$ is independent of $(X_{i})_{i\in[n]}$, $W_{1}$, $W_{2}$ and $W_{3}$ are independent. Moreover, $W_{1}$ is an exponential random variable of parameter $(k-\ell)\lambda$, $W_{2}$ is an exponential random variable of parameter $(m-\ell)\lambda$ and $W_{3}$ is an exponential random variable of parameter $\ell\lambda$. Then, $V_{1}=\min(W_{1},W_{3})=\min_{i\in S}X_{i}$ and $V_{2}=\min(W_{2},W_{3})=\min_{i\in T}X_{i}$. By Corollary 1.8, we have that
$R(V_{1},V_{2})\leq R((S,X_{S}),(T,X_{T}))=\frac{\ell}{\sqrt{mk}}.$
Hence, (54) holds with $\lambda_{1}=(k-\ell)\lambda$, $\lambda_{2}=(m-\ell)\lambda$ and $\lambda_{3}=\ell\lambda$. Therefore, (54) holds for rational $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. For general $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$, take $\lambda^{(N)}_{i}=[N\lambda_{i}]/N$ for $i=1,2,3$ and $N\geq 1$. For each $N\geq 1$, the corresponding random variables are $W^{(N)}_{1},W^{(N)}_{2},W^{(N)}_{3},V^{(N)}_{1}$ and $V^{(N)}_{2}$. As $N\to\infty$, $(V^{(N)}_{1},V^{(N)}_{2})$ converges in distribution to $(V_{1},V_{2})$. By Lemma 2.3, we have that
@@
By similar arguments, we actually have the following upper bound:
@@
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. (real) random variables. Let $1\leq\ell\leq m\leq n$. Then, we have that
@@ 6 Open problems
    Let $(X_{t},Y_{t})_{t\in[0,1]}$ be a two-dimensional Lévy bridge. Is there an expression for the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$?
@@
    By [BS25, Corollary 2.2], the upper bound in (55) is sharp if $X_{i}$ follows exponential distribution. Is it also sharp for other distributions?

### v3

@@ 5.6 Examples
Firstly, we consider an example studied in [BS25]. They calculated the maximal correlation for the bivariate Marshall-Olkin exponential distribution. We state their result in the following theorem:
@@
Let $W_{1},W_{2}$ and $W_{3}$ be independent exponential random variables. The parameter of $W_{i}$ is $\lambda_{i}>0$ for $i=1,2,3$. Let $V_{1}=\min(W_{1},W_{3})$ and $V_{2}=\min(W_{2},W_{3})$. Then we have that
@@
using Corollary 1.7 and Lemma 2.3.
Let $1\leq\ell+1\leq m\leq n$. Let us take independent random variables $X_{1},X_{2},\ldots,X_{n}$. Each $X_{i}$ is an exponential random variable with parameter $\lambda$. Let
$W_{1}=\min_{i=1,2,\ldots,\ell}X_{i},W_{2}=\min_{i=m+1,m+2,\ldots,n}X_{i}\text{ and }W_{3}=\min_{i=\ell+1,\ell+2,\ldots,m}X_{i}.$
Then $W_{1}$, $W_{2}$ and $W_{3}$ are independent. Moreover, $W_{1}$ is an exponential random variable with parameter $\ell\lambda$, $W_{2}$ is an exponential random variable with parameter $(n-m)\lambda$, and $W_{3}$ is an exponential random variable with parameter $(m-\ell)\lambda$. Let $V_{1}=\min(W_{1},W_{3})$ and $V_{2}=\min(W_{2},W_{3})$. By Corollary 1.7, we have that
$R(V_{1},V_{2})\leq R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{j=\ell+1}^{n}\delta_{X_{j}}\right)=\frac{m-\ell}{\sqrt{m(n-\ell)}}.$
Hence, (54) holds with $\lambda_{1}=\ell\lambda$, $\lambda_{2}=(n-m)\lambda$ and $\lambda_{3}=(m-\ell)\lambda$. Therefore, (54) holds for rational $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. For general $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$, take $\lambda^{(N)}_{i}=[N\lambda_{i}]/N$ for $i=1,2,3$ and $N\geq 1$. For each $N\geq 1$, the corresponding random variables are $W^{(N)}_{1},W^{(N)}_{2},W^{(N)}_{3},V^{(N)}_{1}$ and $V^{(N)}_{2}$. As $N\to\infty$, $(V^{(N)}_{1},V^{(N)}_{2})$ converges in distribution to $(V_{1},V_{2})$. By Lemma 2.3, we have that
@@
Using similar arguments, we obtain the following upper bound:
@@
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. real-valued random variables. Let $1\leq\ell+1\leq m\leq n$. Then we have that
@@ 6 Open problems
    Let $(X_{t},Y_{t})_{t\in[0,1]}$ be a two-dimensional Lévy bridge. Is there an expression for the maximal correlation coefficient $R((X_{t})_{t\in[0,1]},(Y_{t})_{t\in[0,1]})$?
@@
    By [BS25, Corollary 2.2], the upper bound in (55) is sharp if $X_{i}$ follows an exponential distribution. Is it also sharp for other distributions?
Appendix A Proof of Lemma 2.1
Lemma 2.1 is derived from the fact that the operator norm of the composition of two operators is not greater than the product of the operator norms of these two operators. In the proof, the Markov property of $(X,Y,Z)$ is crucially used.
Proof of Lemma 2.1.
${L^{2}_{0}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{3}}$${L^{2}_{0}(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{4}}$$\scriptstyle{\iota}$${L^{2}_{0}(Y,Z)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{5}}$${L^{2}_{0}(Z)}$ (56)
where $\pi_{1},\pi_{2},\pi_{3},\pi_{4}$ and $\pi_{5}$ are orthogonal projections and $\iota$ is the injection. Specifically, $\pi_{1}:W\mapsto E(W|Z)$, $\pi_{2}:W\mapsto E(W|Y)$, $\pi_{3}:W\mapsto E(W|Y,Z)$, $\pi_{4}:W\mapsto E(W|Z)$ and $\pi_{5}:W\mapsto E(W|Z)$ are conditional expectation operators. Hence, we have that
where $\|\cdot\|$ denotes the operator norm. ∎
Appendix B Proof of Lemma 2.3
To prove Lemma 2.3, we need to use the classical result that $L^{p}$ functions can be approximated by bounded continuous functions. We give a precise statement as follows:
Lemma B.1.
Consider a metric space S with the Borel $\sigma$-field $\mathcal{S}$, a bounded measure $\mu$ on $(S,\mathcal{S})$ and a constant $p>0$. Then the bounded continuous functions on S are dense in $L^{p}(S,\mathcal{S},\mu)$. Thus, for any $f\in L^{p},$ there exist bounded continuous functions $f_{1},f_{2},\ldots:S\to\mathbb{R}$ with $\|f_{n}-f\|_{p}\to 0$.
The above lemma is precisely Lemma 1.37 in [Kal21].
Proof of Lemma 2.3.
Let $(X_{n},Y_{n})$ be a sequence of random variables taking values in the measurable space $(S,\mathcal{S})$. Suppose that $(X_{n},Y_{n})$ converges weakly to $(X,Y)$ as $n\to\infty$. It suffices to show that
$R(X,Y)\leq\liminf_{n\to\infty}R(X_{n},Y_{n}).$
For any $\varepsilon>0$, by the definition of the maximal correlation coefficient, there exist $\varphi$ and $\psi$ such that
$R(X,Y)\leq\rho(\varphi(X),\psi(Y))+\varepsilon$
and that $\operatorname{Var}(\varphi(X))>0$ and $\operatorname{Var}(\psi(Y))>0$. By Lemma B.1 with $p=2$, for any $\delta>0$, there exist bounded continuous functions $f$ and $g$ such that
$E[(f(X)-\varphi(X))^{2}]<\delta,E[(g(Y)-\psi(Y))^{2}]<\delta.$
$\lim_{n\to\infty}\rho(f(X_{n}),g(Y_{n}))=\rho(f(X),g(Y)).$
By the definition of maximal correlation coefficients,
$\rho(f(X_{n}),g(Y_{n}))\leq R(X_{n},Y_{n}).$
Hence, for any $\varepsilon>0$, we have that
By taking $\varepsilon\to 0$, we obtain the desired result. ∎
Appendix C Proof of Remark 1.2
In this part, we prove Remark 1.2.
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. real-valued random variables. For fixed $m\in[n]$, let $T$ be a uniform random subset of $\{1,2,\ldots,n\}$ with size $m$. For $i=1,2,\ldots,n$, define
where $\partial$ is a special point outside of $\mathbb{R}$. By the Madiman-Barron inequality, for $n\geq 1$,
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{m/n}.$
Then $(\sum_{i=1}^{n}X_{i},\sum_{i=1}^{n}Y_{i}1_{Y_{i}\neq\partial})$ has the same joint distribution as $(S_{n},S_{m})$ in (4). Hence, we see that
Analogous results also hold for i.i.d. random vectors $X_{1},X_{2},\ldots,X_{n}$.
Acknowledgement
The authors would like to thank the anonymous referees for their careful reading and valuable suggestions. We are grateful to Professor Renming Song for his careful reading and helpful suggestions to improve the English writing. Finally, we are grateful to Professor Pei Xu who drew our attention to the problem of maximal correlations.